OpenStudy (anonymous):

A gardener wants to grow carrots along the side of her house. To protect the carrots from wild rabbits, the plot must be enclosed by a wire fence. The gardener wants to use 20 feet of fence material left over from a previous project. Assuming that she constructs a rectangular plot, using the side of her house as one edge, estimate the area of the largest plot she can construct. ___________square feet.

OpenStudy (anonymous):

Thank you for taking the time to drop by and helping me out:)

OpenStudy (anonymous):

This is an optimization problem. Amount of fence used= 2w+l=20. A=lw. A=(20-2w)w. A'=w(-2)+20-2w. Set A' to zero in order to optimize the function. 0=-4w+20 -> w=5 -> when w= 5 l must equal 10 as 2w+l=20

OpenStudy (anonymous):

good?

OpenStudy (anonymous):

hmmm im still working on it...

OpenStudy (anonymous):

the key to optimization problems is that you almost always need to substitute something.

OpenStudy (anonymous):

hmmm so I would use ...?

OpenStudy (anonymous):

oh wait you already solved it didn't you heheh...hold up let me check it on wiley plus this math system ...

OpenStudy (anonymous):

so were substituting -- the constraint of this function -- 2w+l = 20.

OpenStudy (anonymous):

I tried 10 for the answer and it didnt work >.< hehe i must have done something wrong

OpenStudy (anonymous):

Thank you!!! it was 50 right:) lol I just got back to earth, thanks...

OpenStudy (anonymous):

booyah first answer, first medal. glad to help you out.

OpenStudy (anonymous):

Nice, first medal:) YW.